### v1

# Optimal sensing of photon addition and subtraction on nonclassical light

Abstract
We demonstrate that the Wasserstein distance $W_{1}$ corresponding to optical tomograms of nonclassical states faithfully captures changes that arise due to photon addition to, or subtraction from, these states. $W_{1}$ is a true measure of distance in the quantum state space, and is sensitive to the underlying interference structures that arise in the tomogram after changes in the photon number. Our procedure is universally applicable to the cat and squeezed states, the former displaying the characteristic negativity in its Wigner function, while the latter does not do so. We explicate this in the case of the squeezed vacuum and even coherent states and show that photon addition (or subtraction) is mirrored in the shift in the intensity of specific regions in the tomogram. Further, we examine the dependence of $W_{1}$ on the squeezing parameter, and its sensitivity to different quadratures.
I Introduction
An important aspect of quantum metrology is to increase the extent of sensing of nonclassical properties of quantum states beyond the standard quantum limit [1]. The strongly squeezed vacuum state sent through a Mach-Zehnder interferometer [2], the two-mode squeezed vacuum state [3], and the photon added two-mode squeezed vacuum state [4] are examples of ideal candidates for this purpose. It has been shown that photon addition to the squeezed vacuum state (SVS) and the even coherent state (ECS) leads to a shift in the phase space interference features in the corresponding Wigner function as well as modifications in the associated sub-Planck structures, leading to significant metrological changes [5, 6]. However, state reconstruction is in general an arduous task for continuous variable systems with large dimensional Hilbert spaces. In this paper we describe an optimal program for identifying photon number changes in nonclassical light solely from its tomogram. The tomogram is the starting point in any state reconstruction procedure. While, in principle, the inverse Radon transformation of the tomogram gives the Wigner function, the stepwise procedure to be implemented for this purpose is, in general, challenging, particularly for multipartite systems. As a first step, we demonstrate the efficacy of our procedure both for the SVS and the ECS. The details concerning the ECS are discussed in Sec. III of the supplementary material (SM). The procedure that we outline would possibly be optimal for multipartite states as well.
The SVS and other squeezed states of light are of great current interest, as they are very useful in beating the standard quantum limit [7], in quantum error correction [8, 9], enhancing quantum sensing [10], in the metrology of absorption and gain parameters [11, 12], generation of non-standard types of photon blockade [13] and provide the potential for optical switching between the normal and superradiant phases of light [14]. The SVS has diverse applications in, for instance, photoelectric detection [15], gravitational wave detectors [16, 17], parameter estimation in bilinear Hamiltonians [18]. It is therefore not surprising that optimal methods to generate single-mode squeezed light, [19, 20], and obtain a quick and reliable estimation of its extent of degradation due to decoherence, using machine learning tools [21] are being examined extensively.
Investigations on the effect of the addition of a photon to, or its removal from the light field have been undertaken for decades (see, for instance, [22]). Different aspects pertaining to photon addition to nonclassical light  [23, 24, 25, 26, 27]) have been reported in the literature. Quick generation of heralded optical cat states by addition of photons has been experimentally realised [28]. Investigations on photon subtraction from nonclassical light are also of considerable importance (see, for instance, [29, 30, 31, 32, 33, 25, 34, 35, 36, 37, 38]). In what follows we point out among other results, a simple and elegant procedure to detect photon addition or subtraction to the SVS by observing and quantifying changes in its tomogram. This procedure also holds for other nonclassical states and is more optimal than performing investigations with their Wigner functions, although the latter also display substructures.
An important feature of the tomogram of a quantum state is that it is a set of probability distributions in different quadratures, obeying the laws of classical probability theory. In general, we shall therefore compare the new state (after photon addition or subtraction) with the reference state, quantifying the difference, using the notion of a distance between probability distributions. In what follows we use the Wasserstein distance $W_{1}$ [39] for our purpose. We point out that $W_{1}$ satisfies the properties of a distance metric in contrast to other similar quantifiers such as the Kullback-Leibler divergence [40] and the Bhattacharyya distance [41], which are useful in defining entanglement indicators [42].
Each tomogram is represented as an image or pattern as explained later. (Hence $W_{1}$ is a pattern measure and has extensive applications in image processing and machine learning protocols.) In earlier literature $W_{1}$ has been effectively used to quantify the divergence between different eigenstate of the simple harmonic oscillator, a particle in a one-dimensional box and the state of light at different instants when it interacts with an atomic medium [43]. Here, we demonstrate that quantifying the changes in the tomogram using $W_{1}$ suffices to identify changes in the photon number. The significant advantage, as mentioned earlier, is that state reconstruction can be avoided for this purpose. This exercise also reveals interesting links between the SVS and the ECS from a tomographic point of view, for it augments prior knowledge on ‘Janus-faced’ partners reported in the literature [44, 45] (See Sec. III of SM.)
This paper is organized as follows. In Sec. II we discuss the role played by $W_{1}$ in comparing tomographic patterns. In Sec. III we use $W_{1}$ to distinguish between SVS and the photon added SVS. A similar exercise is carried out in Sec. IV on the effect of photon subtraction to the SVS. We conclude with a brief summary and outlook in Sec. V.
II Tomographic patterns and the Wasserstein distance
Consider a single-mode radiation field with photon creation and annihilation operators $\hat{a}^{\dagger}$ and $\hat{a}$ respectively. The set of rotated quadrature operators [46],
$\mathbb{\hat{X}_{\theta}}=\left(\hat{a}^{\dagger}e^{i\theta}+\hat{a}e^{-i\theta}\right)/\sqrt{2}.$ (1)
Here $\theta$ ($0\leqslant\theta<\pi$) is the phase of the local oscillator in the standard homodyne measurement setup. The $x$-quadrature (resp. $p$-quadrature) corresponds to $\theta=0$ (resp. $\theta=\pi/2$). Equation  (1) constitutes a quorum of observables which carry complete information about the single-mode state with density matrix $\hat{\rho}$. The corresponding optical tomogram $w(X_{\theta},\theta)$ [47] is defined as
$w(X_{\theta},\theta)=\langle X_{\theta},\theta|\hat{\rho}|X_{\theta},\theta\rangle,$ (2)
where
$\mathbb{\hat{X}_{\theta}}|X_{\theta},\theta\rangle=X_{\theta}|X_{\theta},\theta\rangle.$ (3)
For a normalized pure state $|\psi\rangle$ it is evident that the tomogram is given by $|\psi(X_{\theta},\theta)|^{2}$ for a given $\theta$. The experimentally measured quantity is $w(X_{\theta},\theta)$ (i.e., only the diagonal elements of ${\hat{\rho}}$ for a given $\theta$). The completeness relation is given by
[…]
where $\{|X_{\theta},\theta\rangle\}$ forms a complete basis for every $\theta$. For computational purposes it is advantageous to expand $w(X_{\theta},\theta)$ in the Fock basis [48].
The tomogram exhibits the symmetry property
$w(X_{\theta},\theta+\pi)=w(-X_{\theta},\theta),$ (5)
and is commonly represented as a pattern with $X_{\theta}$ as the abscissa, and $\theta$ as the ordinate. Single-mode optical tomograms for the SVS, one-photon added SVS and the two-photon added SVS are shown in Fig. 1. For state reconstruction, although it is sufficient to work with the range $0\leqslant\theta<\pi$, the tomogram plotted for $0\leqslant\theta<2\pi$ helps visualize various features better.
In what follows, we will use the Wasserstein distance to compare two tomograms $w_{{\rm A}}(X_{\theta_{{\rm A}}},\theta_{{\rm A}})$ and $w_{{\rm B}}(X_{\theta_{{\rm B}}},\theta_{{\rm B}})$ corresponding to two distinct single-mode radiation fields A and B. Without loss of generality, we first set $\theta_{A}=\theta_{B}=0$ (the $x$-quadrature for each field). The two normalized one-dimensional probability density functions (PDFs) are represented by $f(x)$ and $g(x)$, corresponding to the fields A and B, respectively. The Wasserstein distance $W_{1}$ essentially quantifies the minimum cost in transporting one PDF to the other optimally. It is given in terms of the corresponding cumulative distribution functions (CDFs) $F(x)$ and $G(x)$ as
[…]
Note that $W_{1}$ is a true distance metric, in the sense that $W_{1}(F,F)=0$, $W_{1}(F,G)=W_{1}(G,F)$ and it satisfies the triangle inequality. We now proceed to obtain $W_{1}$ between the SVS and its photon added and photon subtracted counterparts. Table 1 in the SM lists the states and their relevant probability amplitudes.
III Photon Added Squeezed Vacuum States
The squeezed vacuum $|\xi\rangle={\hat{S}}(\xi)|0\rangle$, where the squeezing operator ${\hat{S}}(\xi)=\exp\left[\frac{1}{2}({\xi}^{*}{\hat{a}}^{2}-\xi{\hat{a}}^{\dagger 2})\right]$. Here $\xi=re^{i\phi}$, $r\geqslant 0$ and $\phi\in[0,2\pi)$. For $\phi=0$ (resp. $\phi=\pi$) squeezing is along the $x$-quadrature (resp. $p$-quadrature). It can be shown that [49]
${\hat{S}}(\xi)=\exp\left[-\frac{1}{2}e^{-i\phi}\tanh(r)~{}{\hat{a}}^{\dagger 2}\right]\left[(\cosh r)^{-1/2}\sum_{n=0}^{\infty}\frac{({\rm sech}~{}r-1)^{n}}{n!}(\hat{a}^{\dagger}\hat{a})^{n}\right]\exp\left[\frac{1}{2}e^{i\phi}\tanh(r)~{}{\hat{a}}^{2}\right].$ (7)
Hence $|\xi\rangle$ can be obtained by the action of an exponential operator which is a function of $a^{\dagger 2}$ and its higher powers alone, i.e., $\exp(g{\hat{a}}^{\dagger 2})|0\rangle$, where $g$ is an appropriate constant. We will draw attention to this feature in Sec. III of SM, where we discuss the ECS.
Figure 1: Left to right: Single-mode optical tomograms corresponding to (a) $|\xi\rangle$, (b) $|\xi,1\rangle$ (or $|\xi,-1\rangle$), (c) $|\xi,2\rangle$, and (d) $|\xi,-2\rangle$. The squeezing parameter $\xi=re^{i\phi}$ ($r=1/\sqrt{2}$ and $\phi=0$). The bright central region at $\theta=0$, $\pi$, $2\pi$ in (a) indicates a peak in the PDF. Addition or subtraction of a single photon to the SVS $|\xi\rangle$ destroys this peak (the central dark region in (b)). With the addition or subtraction of two photons the bright central region reappears ((c) and (d)).
The SVS $|\xi\rangle$ can be expressed in the Fock basis as
$|\xi\rangle=\frac{1}{\sqrt{\cosh r}}\sum_{n=0}^{\infty}(-1)^{n}\frac{\sqrt{(2n)}!}{2^{n}n!}e^{in\phi}\tanh^{n}r~{}|2n\rangle.$ (8)
By suitably normalizing $\hat{a}^{\dagger m}|\xi\rangle$ ($m=1,2,\dots$) we obtain the $m$-photon added SVS. Of direct relevance to us are the $1$-photon added SVS
$|\xi,1\rangle=\mathcal{N}_{1}\sum_{n=0}^{\infty}(-1)^{n}\frac{\sqrt{(2n+1)}!}{2^{n}n!}e^{in\phi}\tanh^{n}r~{}|2n+1\rangle,$ (9)
where $\mathcal{N}_{1}=(\cosh r)^{-3/2}$, and the $2$-photon added SVS
$|\xi,2\rangle=\mathcal{N}_{2}\sum_{n=0}^{\infty}(-1)^{n}\frac{\sqrt{(2n+2)}!}{2^{n}n!}e^{in\phi}\tanh^{n}r~{}|2n+2\rangle,$ (10)
with $\mathcal{N}_{2}=\left(3\cosh^{5}r-\cosh^{3}r\right)^{-1/2}$. The tomogram for the SVS does not display alternating intensity fringes. However, on addition of photons to the SVS, such fringes appear. Depending on the number of photons added there are shifts in the alternating bright and dark intensity fringes seen at $\theta=0$, $\pi$ and $2\pi$, with $\phi=0$ (see Fig. 1 and its zoomed-in version in Fig. 2). The number of dark bands corresponds to the number of added photons (Figs. 1(b) and (c)) or to the number of subtracted photons (Figs. 1(b) and (d)).
Figure 2: Left to right: Zoomed-in single-mode optical tomograms corresponding to (a) $|\xi\rangle$, (b) $|\xi,1\rangle$ (or $|\xi,-1\rangle$), (c) $|\xi,2\rangle$, and (d) $|\xi,-2\rangle$ from $\theta=0$ to $\theta=\pi/16$. The parameter values are as in Fig. 1. The shift in the tomographic intensity pattern of the SVS $|\xi\rangle$ with addition or subtraction of photons shown in Fig. 1 is magnified here for better clarity.
Figure 3: Probability distribution function (PDF) along the $x$-quadrature corresponding to $|\xi\rangle$ (blue), $|\xi,1\rangle$ (orange), and $|\xi,2\rangle$ (green) for $r=$ (a) $0.38$ and (b) $0.52$. The other parameter values are as in Fig. 1. At $x=0$, the PDF for the SVS $|\xi\rangle$ and the two-photon added SVS $|\xi,2\rangle$ are maximum, while the one-photon added SVS $|\xi,1\rangle$ has a corresponding minimum. This is reflected in the shift in the intensity patterns seen in Figs. 1 and 2.
To quantify the extent of the shift in the fringes seen in the tomograms with addition of photons, we now examine the corresponding PDFs along the $x$-quadrature (see Fig. 3). The corresponding CDFs are calculated using the PDFs. We compute $W_{1}$ treating the SVS as the reference state. In Fig. 4 we show $W_{1}$ between the SVS and the $1$-photon added SVS (black circles), and also between the SVS and the $2$-photon added SVS (red asterisks) as functions of the squeezing parameter $r$. The range of $r$ in the figure is chosen to be experimentally relevant. The two distances are equal at $r\approx 0.45$ beyond which there is a crossover, and the trends in $W_{1}$ change. However, this crossover behavior is absent in quadratures other than the direction of squeezing (see Fig. 1 in the SM). The quadrature variance along the direction of squeezing computed for the SVS and the photon added SVS are shown in Fig. 2 in the SM. The exponent in the fall-off as a function of the squeezing parameter in the case of 1 and 2 photon addition to the SVS is different from that of the SVS as expected.
Figure 4: Wasserstein distance $W_{1}$ between $|\xi\rangle$ and $|\xi,1\rangle$ [or $|\xi,-1\rangle$] (black circles), $|\xi\rangle$ and $|\xi,2\rangle$ (red asterisks), and $|\xi\rangle$ and $|\xi,-2\rangle$ (green triangles) as a function of the squeezing parameter $r$, along the $x$-quadrature. The other parameters are as in Fig. 1. The crossover between $W_{1}$ for one photon addition and two photon addition is at $r\approx 0.45$ and that between one photon subtraction (same as one photon addition) and two photon subtraction is at $r\approx 0.66$. For distinguishing between the different states the crossover neighborhood should be avoided and an appropriate range of parameter values chosen to clearly identify the difference between the relevant Wasserstein distances.
IV Photon Subtracted Squeezed Vacuum States
Photon subtraction from the SVS is carried out in theory by repeated application of $\hat{a}$ on the SVS and suitably normalizing the resulting state. The $1$-photon subtracted SVS
$|\xi,-1\rangle=\mathcal{K}_{-1}\sum_{n=0}^{\infty}(-1)^{n}\frac{\sqrt{(2n+1)}!}{2^{n}n!}e^{in\phi}\tanh^{n}r~{}|2n+1\rangle,$ (11)
with $\mathcal{K}_{-1}=-e^{i\phi}(\cosh r)^{-3/2}$. The $2$-photon subtracted SVS
$|\xi,-2\rangle=\mathcal{K}_{-2}\sum_{n=2}^{\infty}\frac{(-1)^{n}}{\sqrt{(2n-2)!}}\frac{(2n)!}{2^{n}n!}e^{in\phi}\tanh^{n}r~{}|2n-2\rangle,$ (12)
[…]
In Fig. 4 we consider the plot of $W_{1}$ between |$\ket{\xi}$ and |$\ket{\xi,-1}$ (equivalently |$\ket{\xi,1}$) versus $r$ (black circles). When compared with $W_{1}$ between |$\ket{\xi}$ and |$\ket{\xi,-2}$ (green triangles), we see that the two distances are equal at $r\approx 0.66$, beyond which there is a crossover, and the trends in $W_{1}$ shift. As in the case of photon addition, such crossovers are not seen in other quadratures (Fig. 4 in the SM).
V Concluding remarks
In continuous variable systems, quantum states could suffer from errors arising through photon loss or gain. A precise estimate of the number of such photons is a prerequisite for error correction. High precision detection is possible with states with large sensitivity or small variance in those quadratures where the changes arise due to such errors [50]. We have shown that, for appropriate ranges of the state parameter, the effect of one or two photon addition can be captured in the Wasserstein distances computed between the nonclassical reference state (the squeezed vacuum, for instance) and its photon added counterparts. Similar results hold in the case of the even cat state over an appropriate range of parameter values. It is worth noting that noise estimation gets enhanced by addition of photons to the original nonclassical state [5, 51].
Our approach is based on examining the qualitative features of the relevant optical tomograms (histograms of experimental data). In particular we have examined the shifts in the intensities of nonclassical light in specific quadratures, arising due to photon addition or subtraction. The PDFs comprising the relevant tomograms corroborate our findings. We have demonstrated the usefulness of this tomographic approach both for identifying qualitative changes and for quantifying such changes using the Wasserstein distance. In particular, we have identified specific small ranges of values of state parameters (the crossovers) which are not sufficiently sensitive to photon addition to, or removal from the nonclassical reference state. Furthermore, we have also obtained a sufficiently wide range of experimentally accessible parameter values where it is possible to quantify with precision the extent of change in the photon number by computing the Wasserstein distance directly from the tomograms.
In conclusion, we have presented an experimentally optimal, tomogram-based measure of quantum state discrimination for sensing subtle differences between nonclassical states. It is based on the Wasserstein distance, a true metrological measure, that discriminates states with addition and subtraction of photons. Its advantages are illustrated through examples of nonclassical states, ranging from squeezed vacuum to cat states and their generalizations. Our approach suggests the possibility of sensing changes in the photon number directly from the optical tomogram, circumventing detailed state reconstruction. This could be potentially useful for identifying such changes in generic multimode states where reconstructing the Wigner function from the tomogram could pose considerable challenges.
Acknowledgements.
We acknowledge partial support through funds from Mphasis to the Centre for Quantum Information, Computing and Communication (CQuICC), Indian Institute of Technology Madras. SL and VB thank the Department of Physics, Indian Institute of Technology Madras for infrastructural support.